# Electromagnetic Flow Control in Hypersonic Rarefied Environment

### v1

@@ Abstract
The Unified Gas-Kinetic Wave-Particle (UGKWP) method, developed for multiscale simulation of partially ionized plasmas, has been extended to unstructured meshes, enabling the modeling of electromagnetic flows around a hemisphere across near-continuum to rarefied regimes. To the best of our knowledge, this work represents the first application of a multiscale plasma solver to this problem. In our approach, neutrals, ions, and electrons are treated as distinct species, with electrons modeled beyond the conventional fluid approximation. The numerical implementation is validated through comparison with reference solutions for neutral hypersonic flow around a sphere, as well as benchmarking against experimental data for a Mach 4.75 pre-ionized argon flow. In both cases, the UGKWP results show good agreement with the reference and experimental data. The findings reveal that rarefied effects play a significant role in the prediction of electromagnetic flow control, underscoring the necessity of multiscale modeling in plasma flow applications.
@@ 1 Introduction
Over subsequent decades, considerable progress was made in theoretical, computational, and experimental aspects of MHD thermal protection. In 1958, Bush provided an analytical solution describing the magnetic field’s effect on shock standoff distance [3], while Ziemer and Bush later provided experimental validation using a Mach 4.5 sphere model [4], marking important early milestones. Cambel further investigated the Hall effect’s detrimental impact on magnetic control efficiency [5] and conducted corresponding experiments using pre-ionized argon gas [6], creating benchmark datasets that remain valuable for numerical validation. Following these foundational studies, advances in computational fluid dynamics enabled high-fidelity simulations of MHD flows [7, 8, 9]. Recent numerical studies have increasingly incorporated complex physical phenomena, such as rarefied gas effects and wall sheath interactions. For example, Fawley [10] utilized Direct Simulation Monte Carlo (DSMC) methods to model conductivity in transitional flow regimes, while Katsurayama [11] applied DSMC to simulate rarefied gas effects, approximating the Lorentz force as a body force rather than explicitly resolving charged particle trajectories. Parent [12] further quantified the impact of plasma sheaths on the overall flow field.
@@
Despite these advancements, most numerical simulations rely on MHD equations, whose applicability diminishes under high-altitude conditions characterized by large Knudsen numbers. In such regimes, the continuum assumption inherent to MHD-based models breaks down [13]. Accurately capturing flow physics across continuum, transitional, and rarefied regimes thus requires a genuinely multiscale numerical approach. Addressing these challenges, recent developments have led to the Unified Gas-Kinetic Wave-Particle (UGKWP) method, which enables robust modeling of multiscale neutral gas flows from the rarefied to continuum regimes [14, 15]. The UGKWP framework has since been systematically extended to a variety of transport phenomena, including plasma [16, 17], granular flows [18, 19], radiative transfer [20, 21], and phonon transport [22], demonstrating its capability to capture non-equilibrium effects. Most recently, UGKWP has been specifically applied to partially ionized plasma flows [23]. In the continuum regime, UGKWP accurately recovers ideal MHD, dissipative MHD, and two-fluid models; as the flow transitions to rarefied conditions, the method seamlessly shifts to a particle-based description, effectively capturing kinetic plasma behavior and recovering Particle-in-Cell (PIC) dynamics in the collisionless limit.
@@
The remainder of this paper is organized as follows: Section 2 provides a concise overview of the UGKWP method for partially ionized plasma. Section 3 presents the numerical test results. Finally, Section 4 offers the main conclusions of this study.

### v2

@@ Abstract
The Unified Gas-Kinetic Wave-Particle (UGKWP) method, developed for the multiscale simulation of partially ionized plasmas, is applied to model electromagnetic flows around a hemisphere spanning regimes from near-continuum to rarefied conditions. To the best of our knowledge, this study presents the first application of a multiscale plasma solver to such a problem. In the formulation, neutrals, ions, and electrons are treated as distinct species. The numerical implementation is validated through comparisons with reference solutions for neutral hypersonic flow around a sphere, as well as with experimental data for a Mach 4.75 pre-ionized argon flow. In both cases, UGKWP results show close agreement with the reference and experimental data. A further comparative study across different Knudsen numbers demonstrates that rarefaction effects weaken the influence of electromagnetic control. These findings highlight the capability of UGKWP in modeling electromagnetic control problems and underscore the significant role of rarefied effects in predicting flow-control behavior, thereby emphasizing the necessity of multiscale modeling in plasma flow applications.
@@ 1 Introduction
Over subsequent decades, considerable progress was made in theoretical, computational, and experimental aspects of MHD thermal protection. In 1958, Bush provided an analytical solution describing the magnetic field’s effect on shock standoff distance [3], while Ziemer and Bush later provided experimental validation using a Mach 4.5 sphere model [4], marking important early milestones. Cambel further investigated the Hall effect’s detrimental impact on magnetic control efficiency [5] and conducted corresponding experiments using pre-ionized argon gas [6], creating benchmark datasets that remain valuable for numerical validation. Following these foundational studies, advances in computational fluid dynamics enabled high-fidelity simulations of MHD flows [7, 8, 9]. Recent numerical studies have increasingly incorporated complex physical phenomena, such as rarefied gas effects and all plasma sheath and wall interactions. For example, Fawley [10] utilized Direct Simulation Monte Carlo (DSMC) methods to model conductivity in transitional flow regimes, while Katsurayama [11] applied DSMC to simulate rarefied gas effects, approximating the Lorentz force as a body force rather than explicitly resolving charged particle trajectories. Parent [12] further quantified the impact of plasma sheaths on the overall flow field.
@@
Despite these advancements, most numerical simulations rely on MHD equations, whose applicability diminishes under high-altitude conditions characterized by large Knudsen numbers. In such regimes, the continuum assumption inherent to MHD-based models breaks down [13]. DSMC-based approaches become computationally prohibitive due to the need for excessively dense meshes and small time steps, which are restricted by the mean free path scale [14]. Accurately capturing flow physics across continuum, transitional, and rarefied regimes thus requires a genuinely multiscale numerical approach. Addressing these challenges, recent developments have led to the Unified Gas-Kinetic Wave-Particle (UGKWP) method, which enables robust modeling of multiscale neutral gas flows from the rarefied to continuum regimes [15, 16]. The UGKWP framework has since been systematically extended to a variety of transport phenomena, including plasma [17, 18], granular flows [19, 20], radiative transfer [21, 22], and phonon transport [23], demonstrating its capability to capture non-equilibrium effects. Most recently, UGKWP has been specifically applied to partially ionized plasma flows [24]. In the continuum regime, UGKWP accurately recovers ideal MHD, dissipative MHD, and two-fluid models; as the flow transitions to rarefied conditions, the method seamlessly shifts to a particle-based description, effectively capturing kinetic plasma behavior and recovering Particle-in-Cell (PIC) dynamics in the collisionless limit.
@@
In this work, we apply the UGKWP method to modeling the electromagnetic control problems in the rarefied environments. The remainder of this paper is organized as follows: Section 2 provides a concise overview of the UGKWP method for partially ionized plasma. Section 3 presents the numerical test results. Finally, Section 4 offers the main conclusions of this study.
@@ 2 UGKWP Method
2.1 Governing equations
In this subsection, we introduce the physical models and governing equations utilized in this study. Theoretically, the BGK-Maxwell system can be used to model the partially-ionized plasma [24],
$\frac{\partial f_{\alpha}}{\partial t}+\boldsymbol{u}_{\alpha}\cdot\nabla_{\boldsymbol{x}}f_{\alpha}+\frac{q_{\alpha}(\boldsymbol{E}+\boldsymbol{u}_{\alpha}\times\boldsymbol{B})}{m_{\alpha}}\cdot\nabla_{\boldsymbol{u}}f_{\alpha}=Q_{\alpha},$
$\frac{\partial\boldsymbol{B}}{\partial t}+\nabla_{\boldsymbol{x}}\times\boldsymbol{E}=0,\quad\frac{\partial\boldsymbol{E}}{\partial t}-c^{2}\nabla_{\boldsymbol{x}}\times\boldsymbol{B}=-\frac{1}{\epsilon_{0}}\boldsymbol{J},$
$\nabla_{\boldsymbol{x}}\cdot\boldsymbol{E}=\frac{\rho_{c}}{\epsilon_{0}},\quad\nabla_{\boldsymbol{x}}\cdot\boldsymbol{B}=0,$
where $f_{\alpha}=f_{\alpha}(t,\boldsymbol{x},\boldsymbol{u})$ is the microscopic distribution function for species $\alpha$ ($\alpha=i$ for ion and $\alpha=e$ for electron, $\alpha=n$ for atom) at space and time $(\boldsymbol{x},t)$ and microscopic translational velocity $\boldsymbol{u}$. $q_{\alpha}$ is the charge carried by species $\alpha$. $\boldsymbol{E},\boldsymbol{B}$ is electric and magnetic field strength. $\epsilon_{0}$ is the permittivity of free space and $c$ is the light speed, $\boldsymbol{J}$ is the electric current density. $\rho_{c}$ is the electric charge density. $Q_{\alpha}=\sum_{k\neq\alpha}Q_{\alpha k}(f_{\alpha},f_{k})$ is the collision operator of species $\alpha$ between species $k$, approximated by Andries-Aoki-Perthame (AAP) model [25],
$Q_{\alpha}=\frac{g_{\alpha}^{M}-f_{\alpha}}{\tau_{\alpha}},$
where $g_{\alpha}^{M}$ is a Maxwellian distribution,
$g_{\alpha}^{M}=\rho_{\alpha}\left(\frac{m_{\alpha}}{2\pi k\bar{T}_{\alpha}}\right)^{3/2}\exp\left(-\frac{m_{\alpha}}{2k_{B}\bar{T}_{\alpha}}\left(\boldsymbol{u}_{\alpha}-\bar{\boldsymbol{U}}_{\alpha}\right)^{2}\right),$
and post-collision velocity $\bar{\boldsymbol{U}}_{\alpha}$ and temperature $\bar{T}_{\alpha}$ are given as:
$\bar{\boldsymbol{U}}_{\alpha}$ $=\boldsymbol{U}_{\alpha}+\frac{\tau_{\alpha}}{m_{\alpha}}\sum_{k=1}^{N}2\mu_{\alpha k}\chi_{\alpha k}n_{k}\left(\boldsymbol{U}_{k}-\boldsymbol{U}_{\alpha}\right),$
$\bar{T}_{\alpha}$ $=T_{\alpha}-\frac{m_{\alpha}}{3k_{B}}\left(\bar{\boldsymbol{U}}_{\alpha}-\boldsymbol{U}_{\alpha}\right)^{2}+\tau_{\alpha}\sum_{k=1}^{N}\frac{4\mu_{\alpha k}\chi_{\alpha k}n_{k}}{m_{\alpha}+m_{k}}\left(T_{k}-T_{\alpha}+\frac{m_{k}}{3k_{B}}\left(\boldsymbol{U}_{k}-\boldsymbol{U}_{\alpha}\right)^{2}\right),$
where $\mu_{\alpha k}=m_{\alpha}m_{k}/(m_{\alpha}+m_{k})$ is reduced mass, $N$ is number of species in the system. Species $\alpha$’s mean relaxation time $\tau_{\alpha}$ is determined by $1/\tau_{\alpha}=\sum_{k=1}^{m}\chi_{\alpha k}n_{k}$, and interaction coefficient $\chi_{\alpha k}$ for variable hard sphere (VHS) model is [26]:
$\chi_{\alpha k}=\frac{(5-2\omega)(7-2\omega)\sqrt{\pi}}{15}\left(\frac{2k_{B}T_{\alpha}}{m_{\alpha}}+\frac{2k_{B}T_{k}}{m_{k}}\right)^{1/2}\left(\frac{d_{\alpha}+d_{k}}{2}\right)^{2},$
where $d_{\alpha},d_{k}$ are the diameters of the particles. The temperature modifies the molecular diameter as $({d}/{d_{ref}})^{2}=\left({T}/{T_{ref}}\right)^{\frac{1}{2}-\omega},$ where $\omega$ is the power coefficient between viscosity and temperature [27].
For hypersonic applications, the BGK-Maxwell system can be simplified. In many such applications, the ionized gas within the shock layer exhibits low electrical conductivity $\sigma^{cond}$ [7]. This leads to a small induced current density $\boldsymbol{J}$ according to the Ohm’s law $\boldsymbol{J}=\sigma^{cond}(\boldsymbol{E}+\boldsymbol{U}\times\boldsymbol{B})$. Consequently, the magnetic field produced by this current is significantly weaker than the externally applied magnetic field $\boldsymbol{B}_{ext}$. This allows for the neglect of magnetic field distortion by the flow field, enabling the assumption of a static magnetic field for the present investigation. Faraday’s law can be neglected since the static magnetic field will generate no transverse electric field. Furthermore, the hyperbolic divergence cleaning technique realizes the charge neutrality [28]:
$\epsilon_{0}\frac{\partial\boldsymbol{E}}{\partial t}+c^{2}\nabla\phi=0,\quad\frac{\partial\phi}{\partial t}+\nabla\cdot\boldsymbol{E}=\frac{\rho_{c}}{\epsilon_{0}},$
where $\phi$ is the artificial correction potential. An approximate Debye length of $10^{-4}$ m is implemented to alleviate numerical stiffness. Therefore, the final form of the physical model is given as,
$\frac{\partial f_{\alpha}}{\partial t}+\boldsymbol{u}_{\alpha}\cdot\nabla_{\boldsymbol{x}}f_{\alpha}+\frac{q_{\alpha}(\boldsymbol{E}+\boldsymbol{u}_{\alpha}\times\boldsymbol{B}_{ext})}{m_{\alpha}}\cdot\nabla_{\boldsymbol{u}}f_{\alpha}=Q_{\alpha},$ (1)
$\epsilon_{0}\frac{\partial\boldsymbol{E}}{\partial t}+c^{2}\nabla\phi=-\boldsymbol{J},\quad\frac{\partial\phi}{\partial t}+\nabla\cdot\boldsymbol{E}=\frac{\rho_{c}}{\epsilon_{0}}.$
In the fluid limit $\tau_{\alpha}\rightarrow 0$, Eq.(1) turns to a multi-fluid equation as
$\partial_{t}\rho_{n}+\nabla_{\boldsymbol{x}}\cdot\left(\rho_{n}\boldsymbol{U}_{n}\right)=0,$
$\partial_{t}\left(\rho_{n}\boldsymbol{U}_{n}\right)+\nabla_{\boldsymbol{x}}\cdot\left(\rho_{n}\boldsymbol{U}_{n}\boldsymbol{U}_{n}+p_{n}\mathbb{I}\right)=S_{n},$
$\partial_{t}\mathscr{E}_{n}+\nabla_{\boldsymbol{x}}\cdot\left(\left(\mathscr{E}_{n}+p_{n}\right)\boldsymbol{U}_{n}\right)=Q_{n},$
$\partial_{t}\rho_{i}+\nabla_{\boldsymbol{x}}\cdot\left(\rho_{i}\boldsymbol{U}_{i}\right)=0,$
$\partial_{t}\left(\rho_{i}\boldsymbol{U}_{i}\right)+\nabla_{\boldsymbol{x}}\cdot\left(\rho_{i}\boldsymbol{U}_{i}\boldsymbol{U}_{i}+p_{i}\mathbb{I}\right)=q_{i}n_{i}\left(\boldsymbol{E}+\boldsymbol{U}_{i}\times\boldsymbol{B}_{ext}\right)+S_{i},$
$\partial_{t}\mathscr{E}_{i}+\nabla_{\boldsymbol{x}}\cdot\left(\left(\mathscr{E}_{i}+p_{i}\right)\boldsymbol{U}_{i}\right)=q_{i}n_{i}\boldsymbol{U}_{i}\cdot\boldsymbol{E}+Q_{i},$
$\partial_{t}\rho_{e}+\nabla_{\boldsymbol{x}}\cdot\left(\rho_{e}\boldsymbol{U}_{e}\right)=0,$
$\partial_{t}\left(\rho_{e}\boldsymbol{U}_{e}\right)+\nabla_{\boldsymbol{x}}\cdot\left(\rho_{e}\boldsymbol{U}_{e}\boldsymbol{U}_{e}+p_{e}\mathbb{I}\right)=q_{e}n_{e}\left(\boldsymbol{E}+\boldsymbol{U}_{e}\times\boldsymbol{B}_{ext}\right)+S_{e},$
$\partial_{t}\mathscr{E}_{e}+\nabla_{\boldsymbol{x}}\cdot\left(\left(\mathscr{E}_{e}+p_{e}\right)\boldsymbol{U}_{e}\right)=q_{e}n_{e}\boldsymbol{U}_{e}\cdot\boldsymbol{E}+Q_{e},$
$\epsilon_{0}\frac{\partial\boldsymbol{E}}{\partial t}+c^{2}\nabla\phi=-\boldsymbol{J},\quad\frac{\partial\phi}{\partial t}+\nabla\cdot\boldsymbol{E}=\frac{\rho_{c}}{\epsilon_{0}},$
where $\rho_{\alpha},\boldsymbol{U}_{\alpha},\mathscr{E}_{\alpha}$ represent the mass density, macroscopic velocity, and energy density of species $\alpha$, and $p_{\alpha}$ is the thermal pressure of species $\alpha$. $S_{\alpha},Q_{\alpha}$ are the source terms due to cross-species momentum and energy transfer. In the limit of a strong magnetic field, negligible electron mass, and the non-relativistic regime, Eq.(1) simplifies to MHD equations. A detailed asymptotic analysis of this reduction is provided in [24].
2.2 UGKWP method
In this subsection, the UGKWP method to solve the system (1) as it transitions from the rarefied regime to the continuum regime is introduced. First, the multispecies collision operator is rewritten as
$Q_{\alpha}=\frac{g_{\alpha}^{eq}-f_{\alpha}}{\tau_{\alpha}}+\frac{g_{\alpha}^{M}-g_{\alpha}^{eq}}{\tau_{\alpha}},$ (2)
where $g^{eq}_{\alpha}$ is the Maxwellian distribution by collision within the same species, and $g^{M}_{\alpha}$ is the Maxwellian distribution after collisions between different species. Then the operator splitting method is used to solve system (1) and (2) with timestep $\Delta t$,
$\mathcal{L}_{1}:\frac{\partial f_{\alpha}}{\partial t}+\boldsymbol{u}_{\alpha}\cdot\nabla_{\boldsymbol{x}}f_{\alpha}=\frac{g_{\alpha}^{eq}-f_{\alpha}}{\tau_{\alpha}},$ (3)
$\mathcal{L}_{2}:\frac{\partial f_{\alpha}}{\partial t}+\frac{q_{\alpha}(\boldsymbol{E}+\boldsymbol{u}_{\alpha}\times\boldsymbol{B}_{ext})}{m_{\alpha}}\cdot\nabla_{\boldsymbol{u}}f_{\alpha}=0,\quad\epsilon_{0}\frac{\partial\boldsymbol{E}}{\partial t}=-\boldsymbol{J},$ (4)
$\mathcal{L}_{3}:\frac{\partial f_{\alpha}}{\partial t}=\frac{g_{\alpha}^{M}-g_{\alpha}^{eq}}{\tau_{\alpha}},$ (5)
$\mathcal{L}_{4}:\epsilon_{0}\frac{\partial\boldsymbol{E}}{\partial t}+c^{2}\nabla\phi=0,\quad\frac{\partial\phi}{\partial t}+\nabla\cdot\boldsymbol{E}=\frac{\rho_{c}}{\epsilon_{0}}.$ (6)
The variables after each update are denoted as
$\mathcal{L}_{1}:f^{n}_{\alpha}\rightarrow f^{*}_{\alpha},\boldsymbol{W}^{n}_{\alpha}\rightarrow\boldsymbol{W}^{*}_{\alpha},\quad\quad\;\;\mathcal{L}_{2}:f^{*}_{\alpha}\rightarrow f^{**}_{\alpha},\boldsymbol{W}^{*}_{\alpha}\rightarrow\boldsymbol{W}^{**}_{\alpha},\boldsymbol{E}^{n}\rightarrow\boldsymbol{E}^{*},$
$\mathcal{L}_{3}:f^{**}_{\alpha}\rightarrow f^{n+1}_{\alpha},\boldsymbol{W}^{**}_{\alpha}\rightarrow\boldsymbol{W}^{n+1}_{\alpha},\;\mathcal{L}_{4}:\boldsymbol{E}^{*}\rightarrow\boldsymbol{E}^{n+1},\phi^{n}\rightarrow\phi^{n+1}.$
where $\boldsymbol{W}_{\alpha}=(\rho_{\alpha},\rho_{\alpha}\boldsymbol{U}_{\alpha},\rho_{\alpha}\mathscr{E}_{\alpha})$ is mass, momentum and energy densities of species $\alpha$.
First, we introduce algorithms for solving $\mathcal{L}_{1}$. For simplicity, the species subscript $\alpha$ is omitted here. BGK equation without the force term is written as
$\frac{\partial f}{\partial t}+\boldsymbol{u}\cdot\nabla_{x}f=\frac{g-f}{\tau},$
the time-dependent solution of the distribution function at the interface can be written as,
$f\left(\boldsymbol{x},\boldsymbol{u},\boldsymbol{\xi},t\right)=\frac{1}{\tau}\int_{0}^{t}g(\boldsymbol{x}^{\prime},\boldsymbol{u},\boldsymbol{\xi},t^{{}^{\prime}})e^{-\left(t-t^{\prime}\right)/\tau}\mathrm{d}t^{\prime}+e^{-t/\tau}f_{0}\left(\boldsymbol{x}-\boldsymbol{u}t\right),$ (7)
where $\tau$ is the local mean relaxation time. $f_{0}$ is the initial gas distribution function at $t=0$, and $g$ is equilibrium distribution along the characteristic line $\boldsymbol{x}^{{}^{\prime}}=\boldsymbol{x}-\boldsymbol{u}t^{{}^{\prime}}$. Expand the equilibrium distribution function by the Taylor series to the second-order accuracy,
$g^{{}^{\prime}}=g+{\boldsymbol{g}}_{x}\cdot(\boldsymbol{x}^{{}^{\prime}}-\boldsymbol{x})+g_{t}(t^{{}^{\prime}}-t),$ (8)
where $g\equiv g(\boldsymbol{x},\boldsymbol{u},t),g^{{}^{\prime}}\equiv g(\boldsymbol{x}^{\prime},\boldsymbol{u},t^{{}^{\prime}})$. Substitute them into Eq.(7), the numerical multiscale evolution solution for simulation particle can be obtained,
$f(\boldsymbol{x},\boldsymbol{u},t)=\left(1-e^{-t/\tau}\right)g^{+}(\boldsymbol{x},\boldsymbol{u},t)+e^{-t/\tau}f_{0}\left(\boldsymbol{x}-\boldsymbol{u}t\right),$ (9)
where,
$g^{+}\left(\boldsymbol{x},\boldsymbol{u},t\right)=g\left(\boldsymbol{x},\boldsymbol{u},t\right)+\left(\frac{te^{-t\text{/}\tau}}{1-e^{-t\text{/}\tau}}-\tau\right)\boldsymbol{u}\cdot{\boldsymbol{g}}_{x}\left(\boldsymbol{x},\boldsymbol{u},t\right)+\left(\frac{t}{1-e^{-t\text{/}\tau}}-\tau\right)g_{t}\left(\boldsymbol{x},\boldsymbol{u},t\right).$
The equation above describes the solution to the BGK equation, where the distribution function $f$ at time $t$ is a combination of the initial distribution function $f_{0}$ and the Taylor expansion of the equilibrium state $g$. From the particles’ perspective, this equation implies that a particle has a probability of $e^{-t/\tau}$ to freely stream during the time interval $[0,t]$, and a probability of $(1-e^{-t/\tau})$ to collide with other particles. After multiple collisions, the particle distribution will reach the equilibrium state $g^{+}$, whose degree of freedom can be degenerated.
Based on this physical picture, the UGKWP method represents the velocity distribution function in a micro-macro coupled way. A portion of the distribution is captured analytically through the equilibrium distribution function $g^{+}$, while the remaining part is represented by stochastic simulation particles $P_{k}=(m_{k},\boldsymbol{x}_{k},\boldsymbol{u}_{k})$. Here, $m_{k}$ denotes the mass of simulation particle $P_{k}$, which corresponds to a cluster of real gas particles of the same species, and $\boldsymbol{x}_{k}$ and $\boldsymbol{u}_{k}$ represent the position and velocity of the simulation particle $P_{k}$, respectively.
According to Eq.(9), the cumulative distribution function of the particle’s free streaming time $t_{f}$ before the collision is given as
$F\left(t_{f}<t\right)=\exp\left(-t\text{/}\tau\right),$
from which the free stream time $t_{f}$ can be sampled as $t_{f}=-\tau\ln(\eta)$ with $\eta$ a random varible subject to the uniform distribution $\eta\sim U(0,1)$ . For a time step $\Delta t$ , the particles with $t_{f}\geq\Delta t$ will undergo collisionless free streaming, and the particles with $t_{f}<\Delta t$ will experience collisional interactions. The procedure of updating particles in the UGKWP method is
  1. Step 1:
    During the time step, stream each particle $P_{k}$ for a time period of $\min(\Delta t,t_{f,k})$. Then identify and retain the collisionless particles, while removing the collisional particles. Calculate the free-transport flux across cell interfaces contributed by the particles and accumulate the total conservative quantities of the particles $\boldsymbol{W}_{i}^{p}$;
  2. Step 2:
    After updating the macroscopic conservative variables, calculate the total conservative quantities of the collisional particles $\boldsymbol{W}_{i}^{h}$ from the updated conservative quantities $\boldsymbol{W}_{i}$ as $\boldsymbol{W}_{i}^{h}=\boldsymbol{W}_{i}-\boldsymbol{W}_{i}^{p}$;
  3. Step 3:
    At the end of the time step, rebuild the velocity distribution. Calculate the analytical distribution $g^{+,c}$ and resample the collisionless particles from the distribution $g^{+,f}$ according to the updated conservative quantities, sample the free-streaming time $t_{f,k}$ for each particle $P_{k}$ from the cumulative distribution function $F(t_{f}<t)=\exp(-t/\tau)$ $\boldsymbol{W}_{i}^{h}$.
Refer to [24] for a more detailed description of the procedure. In the procedure outlined above, the algorithm for updating the macroscopic conservative variables is introduced in the following paragraph. The evolution of the microscopic velocity distribution is now solved numerically using particles.